OpenStudy (anonymous):

Another proving of identites: tanx/(tan^2x-1) = 1/(tanx-cotx) ? How?

OpenStudy (anonymous):

more algebra

OpenStudy (anonymous):

sir thanks in advance. how it can be solved sir?

OpenStudy (anonymous):

hold on, let me do some damned algebra

OpenStudy (anonymous):

hahah. sorry sir. ok sir.

OpenStudy (anonymous):

ok lets replace,as usual, \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\) in this example there is literally no trig at all! it is nothing but straight up algebra no trig identities after we write everything in terms of sine and cosine

OpenStudy (anonymous):

ok sir.

OpenStudy (anonymous):

making that replacement, i get on the left \[\frac{\frac{b}{a}}{\left(\frac{b}{a}\right)^2-1}\] and stop calling me sir

OpenStudy (anonymous):

wait... no trig identities? wow.

OpenStudy (anonymous):

on the right, i get \[\frac{1}{\frac{b}{a}-\frac{a}{b}}\]

OpenStudy (anonymous):

ok i guess i don't mean actually no trig identities, but the only ones i used are rewriting everything in terms of sine and cosine that is all

OpenStudy (anonymous):

ok ok.

OpenStudy (anonymous):

using basic algebra you can show rather easily that \[\large \frac{\frac{b}{a}}{\left(\frac{b}{a}\right)^2-1}=\frac{1}{\frac{b}{a}-\frac{a}{b}}\] as they are both equal to \[\frac{ab}{b^2-a^2}\]

OpenStudy (anonymous):

for the right hand side, multiply top and bottom by \(ab\) and you get it in one step

OpenStudy (anonymous):

hmmm. (looking a the post)

OpenStudy (anonymous):

for the left hand side, multiply top and bottom by \(a^2\) and you get the same thing

OpenStudy (anonymous):

umm. can't follow, why should I multiply the right side by ab and left side by a^2? where did it came from?

OpenStudy (anonymous):

ok are you good to this step \[\large \frac{\frac{b}{a}}{\left(\frac{b}{a}\right)^2-1}=\frac{1}{\frac{b}{a}-\frac{a}{b}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

replaced all cosines by \(a\) and all sines by \(b\) ok good

OpenStudy (anonymous):

both sides are an ugly mess, i.e. they are both compound fractions, or complex fractions, or whatever they are called too many bars in other words

OpenStudy (anonymous):

yes. true. I honestly don't know what to do next after that.

OpenStudy (anonymous):

so imagine you were in an algebra class and came upon \[\frac{1}{\frac{b}{a}-\frac{a}{b}}\] what would you do?

OpenStudy (anonymous):

reciprocal then cross multiply?

OpenStudy (anonymous):

clear the fractions by multiplying top and bottom by \(ab\)

OpenStudy (anonymous):

i think what you mean really is add the fraction in the denominator and then take the reciprocal that would work too, but it is the donkey way cowboy way is to do this \[\frac{1}{\frac{b}{a}-\frac{a}{b}}\times \frac{ab}{ab}=\frac{ab}{b^2-a^2}\]

OpenStudy (anonymous):

mmm. will the answer becomes ab/(b^2-a^2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now how about this ugly beast \[\large \frac{\frac{b}{a}}{\left(\frac{b}{a}\right)^2-1}\]

OpenStudy (anonymous):

ooohhh. now I get it.

OpenStudy (anonymous):

or if you prefer \[\large \frac{\frac{b}{a}}{\frac{b^2}{a^2}-1}\]

OpenStudy (anonymous):

hahah ugly beast. yeah, why did you come up with a^2?

OpenStudy (anonymous):

now you can see it right? why \(a^2\) will clear the fraction?

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

it is the common denominator of \(a\) and \(a^2\)

OpenStudy (anonymous):

\[\large \frac{\frac{b}{a}}{\frac{b^2}{a^2}-1}\times \frac{a^2}{a^2}=...\] no more complex fraction

OpenStudy (anonymous):

ah ok-ok. so that's how it is done. so the answer is ab/(b^2-a^2) ?

OpenStudy (anonymous):

thanks sir.

OpenStudy (anonymous):

yup as you can see, it was all algebra, no trig at least once you write all in terms of sine and cosine

OpenStudy (anonymous):

imagine trying to do that algebra writing sine and cosine every time would make the confusing algebra almost impossible right?

OpenStudy (anonymous):

thanks for the technique sir. I will use it and solve the 54 identities in the book. :) hahahah. yeah.